if seven can mix 20 drinks in 5 minutes,sue can mix 20 drinks in 10 minutes, jack can mix 20 drinks in 15 minutes,how much time will it take all 3 of them working togehter to mix 20 drinks?
Answer by Alan3354(69443)

if seven can mix 20 drinks in 5 minutes,sue can mix 20 drinks in 10 minutes, jack can mix 20 drinks in 15 minutes,how much time will it take all 3 of them working togehter to mix 20 drinks?
--------------
seven does 4 per minute
Sue does 2 per minute
jack does 4/3 per minute
------------------------- Add
Together 22/3 drinks per minute
= 3/22 minutes per drink
3/22 x 20 = 30/11 minutes